Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night Report 2014-05-05

Observers: Sandra Benitez, Joe Polshaw
Support: Cosimo Inserra
TAT on duty: Maria Teresa Botticella, Emma Walker

--------------------------------------------------------------

WEATHER CONDITIONS:
start: wind 5.4 m/s; temp. 13.8 C; humid. 20%; Thin clouds (mostly on the horizon)
middle: wind 5.6 m/s; temp. 14 C: humid. 12%; Mostly clear, perhaps some very thin cloud cover
end: wind 6 m/s; temp. 13.9; humid. 10%; Thin clouds

The opening of the dome at the beginning of the night was delayed by about 1 hour due to clouds above the telescope. The thick clouds disappeared, but there were thin clouds (cirrus) throughout the night. The seeing was constantly fluctuating between 1" and 1.4" for the first few hours, but settled down to ~0.9" later on, until the end of the night.

--------------------------------------------------------------

INSTRUMENTS: EFOSC2

CALIBRATIONS

Afternoon:
-- EFOSC2 bias, dome flats, arcs

Morning:
-- EFOSC2 arcs

--------------------------------------------------------------


OBSERVATIONS

Target                   Type                         UT-Start         Filter/Grism           Slit (")         Seeing (")        Notes

*** Delay in opening the dome ***

LSQ14mo              Follow (IMAG)           23:59              gri                         N/A              1.2-1.4           (1) 
LTT3864               STD (SPEC)                00:52             Gr11/13/16          1/1.5            0.8-1.2             -
LSQ14bdq             Follow (SPEC)            01:16             Gr13                      1                  0.9-1.3            (2)
LSQ14bdq             Follow (IMAG)           02:32             Ugri                        N/A             1.0-1.4            (3)
LSQ14bip              Class.                       03:19             Gr13                      1                   0.9                  (4)
LSQ14asn             Follow. (SPEC)           03:53             Gr11                       1                  1.0                  (5)
EG274                  STD (SPEC)                05:35             Gr11/13/16            1                  1.0                   -
LSQ14bka             Class.                       06:05             Gr13                       1                 0.8                   (6)
LSQ14bjj               Class.                       07:02             Gr13                       1                 0.8                  (7)
SN2009ip             Follow (SPEC)             07:32             Gr11                       1                 0.9                  (8)
OGLE-079             Follow (SPEC)            09:49             Gr13                       1                 1.0                  (9)


(1) There was a delay in opening the dome due to clouds, so it was too late to do SN2013ai (will attempt it tomorrow). The moon is at 42%, and 43 degrees away, possible thin clouds, and seeing not so good (1.3" at the start) - we can repeat the OB tomorrow night if there is a problem. 4*200s gri
(2) 2*1800s Gr13.
(3) The telescope had to be re-focused, so lost some time. 3*300s U, 2*200s gri
(4) Host z=0.07954 (SDSS). A normal SN type Ia, about 3 weeks after max. z agrees with host.
(5) 2*2700s gr11
(6) Normal SN type Ia, around max., z~0.07
(7) Had to re-focus again... Normal SN type Ia, about a week before max., z~0.08
(8) 4*1800s gr11. Seeing stayed quite stable at 0.9" during the whole OB.
(9) OGLE-2013-SN-079. 1800s gr13 of nearby red source. Into twilight towards the end of the OB. Will repeat tomorrow (hopefully beginning earlier, and even twice if there's enough time).

Comments